OpenStudy (tuxadog):

ON A TIME LIMIT PLZ HELP WILL GIVE MEDAL To make 1 1/2 dozen muffins, a recipe uses 3 1/2 cups of flour. How many cups of flour are needed for every dozen muffins made? Enter your answer as a mixed number in simplest form.

OpenStudy (retireed):

18 / (7/2) = 12 / x Solve for x is the number of cups of flour for one dozen muffins 18x = 12 (7/2) x = 12(7/2) / 18 x = 6(7) /18 x = _____

OpenStudy (retireed):

That's as much as I can provide. All you needed to do is create a ratio of what you know to what you want to solve for.

OpenStudy (tuxadog):

I am confused

OpenStudy (mathmale):

You need to set up a ratio. To make 1 1/2 dozen muffins, you need 3 1/2 cups of flour. This can be written as \[\frac{ 1.5~dozen }{ 3.5~cups }\]

OpenStudy (mathmale):

Equate this ratio to \[\frac{ 1~dozen }{ x }\]

OpenStudy (mathmale):

and then solve for x. Please show your work.

OpenStudy (mathmale):

x will represent the number of cups of flour required to bake 1 dozen muffins.

OpenStudy (tuxadog):

um it is \[1.5\div3.5\]

OpenStudy (tuxadog):

I am so confused

OpenStudy (mathmale):

I outlined for you precisely what you have to do. If my instructions are not clear, please ask questions. See the attached file.

OpenStudy (mathmale):

"Equate" means "set equal to ... "

OpenStudy (mathmale):

Even if you're unsure of what to do, please do something and share your something with me. You need to set up an equation here. On the left side of your equation you should have the fraction (1 1/2 dozen) / (3 1/2 cups). Put in the = sign next. On the right, write the fraction (1 dozen) / x Now solve this equation for x. again, if you're unsure of what to do, ask questions.

OpenStudy (retireed):

Sorry, I can see where I confused you. I had converted a dozen and a half into 18 muffins and obviously 1 dozen equals 12.